3) The vapor pressure of SiC, is 100 mmHg at 5°C and the compound has a normal boiling point of 58 C. What is the enthalpy of vaporization for SiCl? 4) Sodium has a density of 0.971 g/em and crystallizes in a body-centered cubic unit cell. What is the atomic radius of sodium?
Answer #1

3) clasius-clepyron equation

ln(p2/p1) = DH0rxn/R[1/T1 - 1/T2]

P2 = 760 mmhg, T2= 58 c = 331.15 k

P1 = 100 mmhg, T1 = 278.15 k

DHvap = x kj/mol

ln(760/100) = (x/8.314)*((1/278.15)-(1/331.15))

x = DH0vap = 29.3 kj/mol

4) Na-metal = BCC

   a) No of atoms in BCC = 8(corners)*1/8 + 1(center) = 2

   b) mass of Na atom in the cell = 2*23

           = 46 amu

       ( 1 amu = 1.66*10^-24 grams)

                     = 46*1.66*10^-24

                     = 7.636*10^-23 grams


   c) edge length(a) = x cm

          

       volume of unit cell = a^3

                         

       e) density of solid = mass/volume

                         0.971 = (7.636*10^-23)/(a^3)

            a = edge length = 4.284*10^-8 cm

   BCC edgelength(a) = sqrt(16/3)*r = 4.28*10^-8

                    r = radius = 1.85*10^-8 cm